### v1

# Uhlmann and scalar Wilczek-Zee phases of degenerate quantum systems

Abstract
The Wilczek-Zee (WZ) holonomy arises in degenerate states while the Uhlmann holonomy characterizes finite-temperature topology. We investigate possible relationships between the Uhlmann phase and the scalar WZ phase, which reflects the Uhlmann and WZ holonomy respectively, in an exemplary four-level model with two doubly degenerate subspaces. Through exactly solutions, we contrast the behavior of the Uhlmann and WZ connections and their associated phases. In the zero-temperature limit, the Uhlmann phase may or may not agree with the scalar WZ phase of the degenerate ground states due to the underlying topology. Some conditions are derived in an analytic analysis, under which the Uhlmann phase in the zero-temperature limit converges to the scalar WZ phase. We also briefly discuss possible realizations or simulations of the phenomena as well as their implications.
I Introduction
Geometric phases serve as a cornerstone in the study of interesting topological properties of quantum systems. These phases arise from the evolution of a quantum state in the parameter space and reflect the underlying geometric structures, depending only on the path taken rather than the dynamic details of the evolution. The original idea was introduced via the Berry phase [1], which characterizes the phase accumulated by a non-degenerate quantum state during an adiabatic and cyclic evolution. Since its introduction, the Berry phase has become a stepping stone for understanding a wide range of topological phenomena, ranging from the quantum Hall effect to topological insulators [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
For quantum systems with degeneracy, the geometric phase exhibits even richer physics. The Wilczek-Zee (WZ) phase of degenerate systems [14] extends the Berry phase of non-degenerate systems. Unlike the scalar Berry phase, the WZ phase manifests as a unitary matrix acting on the degenerate subspace, reflecting the underlying non-Abelian holonomy. This non-Abelian nature introduces additional complexity and enables the study of topological properties in degenerate quantum systems [15, 16, 17, 18]. Moreover, the WZ phase has found applications across diverse areas, including nuclear rotations of diatomic molecules [19, 20], nuclear quadrupole resonance (NQR) [21, 22], dynamics of deformable bodies [23, 24], molecular Kramers doublets [25], semiconductor heterostructures [26], ion traps [27], etc. More recently, it has also been experimentally observed using atomic Bose-Einstein condensates [28].
While pure states described by wavefunctions are well-suited for studying geometric phases of equilibrium systems at zero temperature, finite-temperature systems are ubiquitous with thermal effects giving rise to mixed states described by density matrices. It is thus crucial to address how geometric phases can be generalized to mixed states. The Uhlmann phase [29, 30] provides a rigorous solution by extending the concept of geometric phase to density matrices through the definition of the parallel-transport condition of purification of density matrices. The Uhlmann phase has been instrumental in revealing topological phase transitions at finite temperatures in various systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Despite its utility, the Uhlmann phase remains less understood than its pure-state counterparts, particularly when it comes to its relationship in the zero-temperature limit to the Berry and WZ phases.
Previous research has revealed interesting examples where the Uhlmann phase reduces, in the zero-temperature limit, to the Berry phase at zero temperature in non-degenerate systems [31, 35, 41], despite their different origins from topologically distinct principal fiber bundles. Those examples motivate a broader examination of how finite-temperature geometric phases relate to possible zero-temperature counterparts, especially in the presence of degeneracy. Here, we will focus on the Uhlmann phase of degenerate states in the zero-temperature limit. To facilitate a fair comparison with the geometric phase from the WZ holonomy of degenerate ground states, we propose to use the scalar Wilczek-Zee phase, which has a similar definition as the Uhlmann phase to reflect the underlying non-Abelian holonomy, instead of the matrix-valued WZ phases in the search of possible relations with the Uhlmann phase in the zero-temperature limit.
To provide a concrete analysis, we study a four-level model with two doubly degenerate subspaces, which readily exhibits a variety of phenomena for contrasting the behavior of the Uhlmann and scalar WZ phases. Moreover, we explicitly derive the Uhlmann connection and the WZ connection that generate the corresponding non-Abelian holonomy. Through two simple examples of the four-level model, we demonstrate some agreements and disagreements by varying the underlying parameter spaces between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. An attempt to analytically understand the zero-temperature limit of the Uhlmann holonomy and phase in contrast to the WZ holonomy and its scalar phase further elucidates some interesting but challenging topological properties behind degenerate quantum systems at and beyond zero temperature.
The rest of the paper is organized as follows. Section II introduces the four-level model with two doubly degenerate subspaces and its properties, defines the scalar WZ phase, and summarizes the explicit expressions for the Uhlmann and WZ connections along with their associated holonomy and phases. Section III examines two specific examples of the four-level model to demonstrate possible relationships between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. Section IV provides a theoretical analysis of the Uhlmann phase in the zero-temperature limit and proposes possible conditions for it to reduce to the scalar WZ phase. Section V discusses some implications of the Uhlmann and scalar WZ phases in possible experiments and simulations. Finally, Section VI summarizes our findings and offers concluding remarks. The Appendix shows some details of the Uhlmann connection.
II Four-level model with two doubly degenerate spaces
II.1 The Hamiltonian
We consider a generic four-level model with two doubly degenerate spaces described by the Hamiltonian
$H=\sum_{i=1}^{5}R_{i}\Gamma^{i}.$ (1)
Here the gamma matrices are defined by $\Gamma^{i}=\sigma_{1}\otimes\sigma_{i}$ for $i=1,2,3$, $\Gamma^{4}=\sigma_{2}\otimes 1_{2}$ and $\Gamma^{5}=\sigma_{3}\otimes 1_{2}$, where $\sigma_{i}$ with $i=1,2,3$ denote the Pauli matrices and $1_{2}$ denotes the $2\times 2$ identity matrix. These gamma matrices satisfy the Clifford algebra $\{\Gamma^{i},\Gamma^{j}\}=2\delta_{ij}1_{4}$ with $1_{4}$ being the $4\times 4$ identity matrix. Throughout the paper, we set $\hbar=1=k_{B}$.
The four eigenstates form two pairs of degenerate states and are given by
$|\psi_{a,c}\rangle$ $=\frac{1}{\sqrt{2R(R\mp R_{5})}}\begin{pmatrix}-R_{3}+\mathrm{i}R_{4}\\ -R_{1}-\mathrm{i}R_{2}\\ R_{5}\mp R\\ 0\end{pmatrix},$
$|\psi_{b,d}\rangle$ $=\frac{1}{\sqrt{2R(R\mp R_{5})}}\begin{pmatrix}-R_{1}+\mathrm{i}R_{2}\\ R_{3}+\mathrm{i}R_{4}\\ 0\\ R_{5}\mp R\end{pmatrix}$ (2)
with $R=\sqrt{\sum_{i=1}^{5}R^{2}_{i}}$. They satisfy the eigenvalue equations
$H|\psi_{a,b}\rangle$ $=+R|\psi_{a,b}\rangle,\quad H|\psi_{c,d}\rangle=-R|\psi_{c,d}\rangle,$ (3)
where each level is twofold degenerate. We introduce the projectors onto the subspaces with eigenvalues ±$\pm R$ as
$P_{+}$ $=|\psi_{a}\rangle\langle\psi_{a}|+|\psi_{b}\rangle\langle\psi_{b}|=\frac{1}{2}\left(1_{4}+\hat{R}_{i}\Gamma^{i}\right),$
$P_{-}$ $=|\psi_{c}\rangle\langle\psi_{c}|+|\psi_{d}\rangle\langle\psi_{d}|=\frac{1}{2}\left(1_{4}-\hat{R}_{i}\Gamma^{i}\right),$ (4)
where $\hat{R}_{i}=R_{i}/R$. Note that each expression for $P_{\pm}$ corresponds to a different representation: the first form is written in the eigen-basis of the Hamiltonian, while the second form is given in terms of the Gamma matrices. In thermal equilibrium, the density matrix is
$\rho=$ $\frac{\mathrm{e}^{-\beta R}}{Z}P_{+}+\frac{\mathrm{e}^{\beta R}}{Z}P_{-}=\frac{1}{4}\left(1_{4}-\tanh(\beta R)\hat{R}_{i}\Gamma^{i}\right),$ (5)
where $\beta=\frac{1}{T}$ is the inverse temperature, and $Z=2\mathrm{e}^{-\beta R}+2\mathrm{e}^{\beta R}=4\cosh(\beta R)$ is the partition function.
II.2 Uhlmann holonomy
In general, we consider a quantum system with a $N$-dimensional Hilbert space. A statistical ensemble is represented by a mixed state, which is described by a density matrix. To define the geometric phase of mixed states, one can introduce the notion of the amplitude through purification of a full-rank density matrix via $W=\sqrt{\rho}\mathcal{U}$, where $\mathcal{U}\in U(N)$ is a unitary phase factor. Conversely, the density matrix can be reconstructed as $\rho=WW^{\dagger}$, which is independent of the choice of the phase factor. For a system in thermal equilibrium described by the canonical ensemble, the density matrix has the diagonal form $\rho=\sum_{n}\lambda_{n}|n\rangle\langle n|$ in the eigen-basis of the Hamiltonian. The purification of the density matrix is then expressed as $W=\sum_{n}\sqrt{\lambda_{n}}|n\rangle\langle n|\mathcal{U}$. One can introduce a pure-state like representation of $W$, called the purified state, as $|W\rangle=\sum_{n}\sqrt{\lambda_{n}}|n\rangle\otimes\mathcal{U}^{T}|n\rangle$. In this way, the Hilbert-Schmidt product introduces an inner product between purified states: $\langle W_{1}|W_{2}\rangle=\text{Tr}(W^{\dagger}_{1}W_{2})$ [42].
For the four-level model described by Eq. (1), we define $\mathbf{R}=(R_{1},R_{2},R_{3},R_{4},R_{5})^{T}$ and consider the system undergoing cyclic evolution along a closed path $C$ described by $\mathbf{R}(t)$ with $0\leq t\leq\tau$ and $\mathbf{R}(0)=\mathbf{R}(\tau)$ in the parameter space. For simplicity, we denote $\rho(t)\equiv\rho(\mathbf{R}(t))$. If the amplitude $W(t)$ satisfies the Uhlmann parallel-transport condition
$W^{\dagger}\dot{W}=\dot{W}^{\dagger}W,$ (6)
then $W(t)$ is called a horizontal lift of $\rho(t)$. The condition further implies the differential equation
$\dot{\mathcal{U}}=-\mathcal{A}_{\text{U}}\mathcal{U},$ (7)
where $\mathcal{A}_{\text{U}}$ is the Uhlmann connection. In general, the curve $W(t)$ is not closed, as the initial and final phase factors may differ by the Uhlmann holonomy, which is obtained from Eq. (7) and has the form
$\mathcal{U}(\tau)=\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}\mathcal{U}(0),$ (8)
where $\mathcal{P}$ denotes path ordering. Using $\rho=\sum_{i}\lambda_{i}|i\rangle\langle i|$, $\mathcal{A}_{\text{U}}$ can be expressed as
$\mathcal{A}_{\text{U}}=-\sum_{ij}|i\rangle\frac{\langle i|[\mathrm{d}\sqrt{\rho},\sqrt{\rho}]|j\rangle}{\lambda_{i}+\lambda_{j}}\langle j|.$ (9)
For four-level model featuring two doubly degenerate subspaces introduced in Eq. (1), a direct evaluation of Eq. (9) shows that the Uhlmann connection simplifies to
$\mathcal{A}_{\text{U}}=-\left[1-\text{sech}\left(\frac{R}{T}\right)\right]\left(P_{+}\mathrm{d}P_{-}+P_{-}\mathrm{d}P_{+}\right)$
$=$ $\frac{1-\text{sech}\left(\frac{R}{T}\right)}{2}\hat{R}_{a}\mathrm{d}\hat{R}_{b}\Gamma^{a}\Gamma^{b}$
$=$ $-\frac{\mathrm{i}}{4}\left[1-\text{sech}\left(\frac{R}{T}\right)\right](\hat{R}_{a}\mathrm{d}\hat{R}_{b}-\hat{R}_{b}\mathrm{d}\hat{R}_{a})\Gamma^{ab},$ (10)
where $\Gamma^{ab}=\frac{\mathrm{i}}{2}[\Gamma^{a},\Gamma^{b}]$. More details are given in Appendix A. One important distinction between Eqs. (9) and (II.2) is that Eq. (9) is given in the eigen-basis of the Hamiltonian, which is also the eigen-basis of the canonical-ensemble density matrix in equilibrium. In contrast, Eq. (II.2) is written in the Gamma-matrix representation with known forms. The latter allows for a direct calculation of the Uhlmann connection in terms of the Gamma matrices:
$\mathcal{A}_{\text{U}}=-\frac{1-\text{sech}\left(\frac{R}{T}\right)}{2R^{2}}M,$ (11)
where the matrices elements of $M$ are given by
$M_{11}$ $=\mathrm{i}(R_{2}\mathrm{d}R_{1}-R_{1}\mathrm{d}R_{2}+R_{4}\mathrm{d}R_{3}-R_{3}\mathrm{d}R_{4}),$
$M_{12}$ $=(R_{1}-\mathrm{i}R_{2})(\mathrm{d}R_{3}-\mathrm{i}\mathrm{d}R_{4})-(R_{3}-\mathrm{i}R_{4})(\mathrm{d}R_{1}-\mathrm{i}\mathrm{d}R_{2}),$
$M_{13}$ $=-R_{5}\mathrm{d}R_{3}+\mathrm{i}R_{5}\mathrm{d}R_{4}+(R_{3}-\mathrm{i}R_{4})\mathrm{d}R_{5},$
$M_{14}$ $=-R_{5}\mathrm{d}R_{1}+\mathrm{i}R_{5}\mathrm{d}R_{2}+(R_{1}-\mathrm{i}R_{2})\mathrm{d}R_{5},$
$M_{21}$ $=(R_{3}+\mathrm{i}R_{4})(\mathrm{d}R_{1}+\mathrm{i}\mathrm{d}R_{2})-(R_{1}+\mathrm{i}R_{2})(\mathrm{d}R_{3}+\mathrm{i}\mathrm{d}R_{4}),$
$M_{22}$ $=\mathrm{i}(-R_{2}\mathrm{d}R_{1}+R_{1}\mathrm{d}R_{2}-R_{4}\mathrm{d}R_{3}+R_{3}\mathrm{d}R_{4}),$
$M_{23}$ $=(R_{1}+\mathrm{i}R_{2})\mathrm{d}R_{5}-R_{5}(\mathrm{d}R_{1}+\mathrm{i}\mathrm{d}R_{2}),$
$M_{24}$ $=R_{5}\mathrm{d}R_{3}+\mathrm{i}R_{5}\mathrm{d}R_{4}-(R_{3}+\mathrm{i}R_{4})\mathrm{d}R_{5},$
$M_{31}$ $=R_{5}\mathrm{d}R_{3}+\mathrm{i}R_{5}\mathrm{d}R_{4}-(R_{3}+\mathrm{i}R_{4})\mathrm{d}R_{5},$
$M_{32}$ $=R_{5}\mathrm{d}R_{1}-\mathrm{i}R_{5}\mathrm{d}R_{2}-(R_{1}-\mathrm{i}R_{2})\mathrm{d}R_{5},$
$M_{33}$ $=\mathrm{i}(R_{2}\mathrm{d}R_{1}-R_{1}\mathrm{d}R_{2}-R_{4}\mathrm{d}R_{3}+R_{3}\mathrm{d}R_{4}),$
$M_{34}$ $=(R_{1}-\mathrm{i}R_{2})(\mathrm{d}R_{3}+\mathrm{i}\mathrm{d}R_{4})-(R_{3}+\mathrm{i}R_{4})(\mathrm{d}R_{1}-\mathrm{i}\mathrm{d}R_{2}),$
$M_{41}$ $=R_{5}\mathrm{d}R_{1}+\mathrm{i}R_{5}\mathrm{d}R_{2}-(R_{1}+\mathrm{i}R_{2})\mathrm{d}R_{5},$
$M_{42}$ $=-R_{5}\mathrm{d}R_{3}+\mathrm{i}R_{5}\mathrm{d}R_{4}+(R_{3}-\mathrm{i}R_{4})\mathrm{d}R_{5},$
$M_{43}$ $=(R_{3}-\mathrm{i}R_{4})(\mathrm{d}R_{1}+\mathrm{i}\mathrm{d}R_{2})-(R_{1}+\mathrm{i}R_{2})(\mathrm{d}R_{3}-\mathrm{i}\mathrm{d}R_{4}),$
$M_{44}$ $=\mathrm{i}(-R_{2}\mathrm{d}R_{1}+R_{1}\mathrm{d}R_{2}+R_{4}\mathrm{d}R_{3}-R_{3}\mathrm{d}R_{4}).$ (12)
At the end of the parallel transport along the loop $C$ in the parameter space, the mixed state acquires the Uhlmann phase
$\theta_{\text{U}}(C)=\arg\langle W(0)|W(\tau)\rangle=\arg\text{Tr}\left[\rho(0)\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}\right].$ (13)
When $T\rightarrow 0$, a system in thermal equilibrium should approach its ground state. Later on, we will analyze the behavior of the Uhlmann connection, Uhlmann holonomy, and Uhlmann phase in the zero-temperature limit in the presence of a degenerate ground-state subspace.
II.3 Wilczek-Zee holonomy
At zero temperature, the ground states are of interest and in equilibrium, they are pure states. Here we briefly summarize the pure-state holonomy in the presence of degenerate states by considering a system evolving adiabatically along a closed path $C$ parameterized by $\mathbf{R}(t)$ for $t\in[0,\tau]$ with $\mathbf{R}(0)=\mathbf{R}(\tau)$. Consequently, the Hamiltonian $H(t)\equiv H(\mathbf{R}(t))$ returns to its initial form at $t=\tau$. Under the cyclic evolution, the pure states in each degenerate subspace undergo a unitary transformation:
$|\psi_{i}(\tau)\rangle$ $=\sum_{j=a,b}\bigl{(}U_{+}\bigr{)}_{ij}|\psi_{j}(0)\rangle,\quad i=a,b,$
$|\psi_{i}(\tau)\rangle$ $=\sum_{j=c,d}\bigl{(}U_{-}\bigr{)}_{ij}|\psi_{j}(0)\rangle,\quad i=c,d.$ (14)
In the ordered basis $\{|\psi_{a}\rangle,|\psi_{b}\rangle,|\psi_{c}\rangle,|\psi_{d}\rangle\}$, the full holonomy, which gives the Wilczek-Zee matrix [14], factorizes into a block-diagonal form:
$U=\begin{pmatrix}U_{-}&0\\ 0&U_{+}\end{pmatrix}=\begin{pmatrix}\mathcal{P}\mathrm{e}^{\mathrm{i}{\int}_{0}^{\tau}\mathbf{A}_{-}\cdot\dot{\mathbf{R}}\mathrm{d}t}&0\\ 0&\mathcal{P}\mathrm{e}^{\mathrm{i}{\int}_{0}^{\tau}\mathbf{A}_{+}\!\cdot\dot{\mathbf{R}}\mathrm{d}t}\end{pmatrix},$ (15)
where the elements of the WZ potentials are given by $\mathbf{A}_{+ij}=\mathrm{i}\langle\psi_{i}|\nabla_{\mathbf{R}}|\psi_{j}\rangle$ for $i,j=a,b$ and $\mathbf{A}_{-ij}=\mathrm{i}\langle\psi_{i}|\nabla_{\mathbf{R}}|\psi_{j}\rangle$ for $i,j=c,d$.
In order to directly compare with the Uhlmann connection later, we define the WZ connection as
$\mathcal{A}_{\pm,ij}=\langle\psi_{i}|\mathrm{d}|\psi_{j}\rangle=-\mathrm{i}\mathbf{A}_{\pm,ij}\cdot\mathrm{d}\mathbf{R}$ (16)
for $i$, $j$=$a$, $b$ or $c$, $d$. This definition differs from the conventional one by a factor of $\mathrm{i}=\sqrt{-1}$, but it does not change any physical essence. Thus, the total WZ connection matrix is
$\mathcal{A}_{\text{WZ}}=\begin{pmatrix}\mathcal{A}_{-}&0\\ 0&\mathcal{A}_{+}\end{pmatrix},$ (17)
and the corresponding WZ holonomy along a loop $C$ is given by $U(C)=\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{WZ}}}$. Unlike the Uhlmann connection given in Eq. (II.2), here the WZ connection is, by definition, expressed in the eigen-basis of the Hamiltonian. To express the WZ connection by the Gamma matrices typically requires a more involved calculation. Nevertheless, we obtain
$\mathcal{A}_{+aa,-cc}$ $=-\mathcal{A}_{+bb,-dd}$
$=\dfrac{R_{2}\mathrm{d}R_{1}-R_{1}\mathrm{d}R_{2}-R_{4}\mathrm{d}R_{3}+R_{3}\mathrm{d}R_{4}}{2\mathrm{i}R(R\mp R_{5})}$ (18)
and
$\mathcal{A}_{+ab,-cd}$ $=\mathcal{A}^{*}_{+ba,-dc}=\frac{R_{1}\mathrm{d}R_{4}-R_{4}\mathrm{d}R_{1}-R_{2}\mathrm{d}R_{3}+R_{3}\mathrm{d}R_{2}+\mathrm{i}(R_{1}\mathrm{d}R_{3}-R_{3}\mathrm{d}R_{1}+R_{2}\mathrm{d}R_{4}-R_{4}\mathrm{d}R_{2})}{2\mathrm{i}R(R\mp R_{5})}.$ (19)
II.4 Scalar Wilczek-Zee phase
Since the explicit forms of $\mathcal{A}_{\text{U}}$ and $\mathcal{A}_{\text{WZ}}$ are available, it is intriguing to check the $T\to 0$ limit of $\mathcal{A}_{\text{U}}$ and see if there is any relationship with $\mathcal{A}_{\text{WZ}}$. A key distinction, however, is that they have been derived in different bases: Eq. (II.2) presents the Uhlmann connection in the Gamma-matrix representation, whereas Eq. (15) gives the WZ connection in the eigen-basis of the Hamiltonian. This apparent discrepancy seems to preclude a direct comparison. Nevertheless, we demonstrate in Appendix A that the Uhlmann connection can also be rewritten in terms of the eigen-basis of the Hamiltonian. Consequently, Eq. (9) can be rewritten as [41]
$\mathcal{A}_{\mathrm{U}}$ $=-\sum_{i,j}\frac{\bigl{(}\sqrt{\lambda_{i}}-\sqrt{\lambda_{j}}\bigr{)}^{2}}{\lambda_{i}+\lambda_{j}}|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|$ (20)
in the eigen-basis of the Hamiltonian. Whenever $|i\rangle$ and $|j\rangle$ lie in the same degenerate subspace, $\lambda_{i}=\lambda_{j}$, causing each diagonal block of $\mathcal{A}_{\mathrm{U}}$ to vanish identically. All nonzero entries of $\mathcal{A}_{\mathrm{U}}$ therefore arise from the coupling between different energy levels. By contrast, $\mathcal{A}_{\mathrm{WZ}}$ is explicitly block-diagonal, with its only nonzero elements confined within each degenerate subspace.
Since an explicit relation between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ is difficult to obtain at present, we shift our focus to the Uhlmann and WZ phases arising from the corresponding holonomy. However, the conventional WZ phase is matrix-valued [14], as shown by the elements in Eq. (15). To enable a direct comparison with the zero-temperature limit of the Uhlmann phase, we introduce the scalar Wilczek-Zee phase associated with the degenerate ground-state subspace:
$\theta_{\text{WZ}}(C)=\arg\text{Tr}_{-}\left(\frac{1}{2}P_{-}U_{-}(C)\right),$ (21)
where $\text{Tr}_{-}$ denotes the trace over the degenerate ground-state subspace, and the prefactor $\frac{1}{2}$ ensures the normalization $\text{Tr}_{-}\left(\frac{1}{2}P_{-}\right)=1$. Although the scalar WZ phase can be similarly defined for the degenerate excited-state subspace, it is irrelevant to the zero-temperature limit of the Uhlmann phase. Therefore, we will focus on the scalar WZ phase of the degenerate ground states in the following. In the absence of degenerate ground states ($D=1$), $\theta_{\text{WZ}}$ reduces to the Berry phase, as expected. When the ground state is $D$-fold degenerate, the normalization factor $\frac{1}{2}$ should be generalized to $\frac{1}{D}$.
To make the comparison between the $T\rightarrow 0$ Uhlmann phase and scalar WZ phase more transparent, we define the analogue of purification of the ground-state projector as $W_{-}=\frac{1}{\sqrt{D}}\sqrt{P_{-}}\mathcal{U}$. In analogy to the Uhlmann phase in Eq. (13), the scalar WZ phase accumulated along a closed path $C(t)$ ($0\leq t\leq\tau$) can then be expressed as
$\theta_{\text{WZ}}(C)=\arg\langle W_{-}(0)|W_{-}(\tau)\rangle.$ (22)
Here the analogies of the purified state $|W_{-}(t)\rangle$ and the inner products are also introduced in a similar fashion. Since the density matrix approaches $\rho\to\frac{1}{D}P_{-}$ as $T\to 0$, the formally similar expressions of the Uhlmann and scalar WZ phases warrant some investigations into their relations. We set out to search for any correspondence between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase in the four-level model described by Eq. (1) because such a correspondence may serve as a bridge to link the topological properties of mixed states with those of degenerate ground states.
III Examples
Here we present two examples of the four-level model discussed in Sec. II to contrast the $T\rightarrow 0$ limit of the Uhlmann phase and the WZ phase of the degenerate ground states.
III.1 A simple illustration
We begin by considering a simple case with $R_{1}=R_{3}=\frac{R}{\sqrt{2}}\sin\theta\cos\phi$, $R_{2}=R_{4}=\frac{R}{\sqrt{2}}\sin\theta\sin\phi$, and $R_{5}=R\cos\theta$ in Eq. (1). Consequently, the parameter manifold is a two-dimensional spherical surface parameterized by $(\theta,\phi)$. After substituting the parameters into Eq. (II.2) and following a straightforward calculation, the Uhlmann connection is given by
$\mathcal{A}_{\text{U}}=$ $-\frac{1-\operatorname{sech}\left(\frac{R}{T}\right)}{2}$
$\times$ $\begin{pmatrix}-\mathrm{i}\sin^{2}\theta\mathrm{d}\phi&0&-\frac{\mathrm{e}^{-\mathrm{i}\phi}}{\sqrt{2}}\Omega&-\frac{\mathrm{e}^{-\mathrm{i}\phi}}{\sqrt{2}}\Omega\\ 0&\mathrm{i}\sin^{2}\theta\mathrm{d}\phi&-\frac{\mathrm{e}^{\mathrm{i}\phi}}{\sqrt{2}}\Omega^{*}&\frac{\mathrm{e}^{\mathrm{i}\phi}}{\sqrt{2}}\Omega^{*}\\ \frac{\mathrm{e}^{\mathrm{i}\phi}}{\sqrt{2}}\Omega^{*}&\frac{\mathrm{e}^{-\mathrm{i}\phi}}{\sqrt{2}}\Omega&0&\mathrm{i}\sin^{2}\theta\mathrm{d}\phi\\ \frac{\mathrm{e}^{\mathrm{i}\phi}}{\sqrt{2}}\Omega^{*}&-\frac{\mathrm{e}^{-\mathrm{i}\phi}}{\sqrt{2}}\Omega&\mathrm{i}\sin^{2}\theta\mathrm{d}\phi&0\end{pmatrix},$ (23)
where $\Omega=\mathrm{d}\theta-\mathrm{i}\cos\theta\sin\theta\mathrm{d}\phi$ is a 1-form, and $\Omega^{*}$ is its complex conjugate. For convenience, we consider a closed path taken along the equator (i.e., $\theta=\frac{\pi}{2}$) of the parameter space, starting from the point $(\theta(0),\phi(0))=(\frac{\pi}{2},0)$. Along the path, the Uhlmann connection simplifies to
$\mathcal{A}_{\text{U}}$ $=\mathrm{i}\frac{1-\operatorname{sech}\left(\frac{R}{T}\right)}{2}\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&0&-1\\ 0&0&-1&0\end{pmatrix}\mathrm{d}\phi$
$=\mathrm{i}\frac{1-\operatorname{sech}\left(\frac{R}{T}\right)}{2}\sigma_{3}\oplus(-\sigma_{1})\mathrm{d}\phi,$ (24)
which is proportional to a constant matrix.
As a consequence, the path-ordering operator $\mathcal{P}$ in Eq. (13) becomes straightforward and may be omitted in the evaluation of the Uhlmann phase. Furthermore, the absence of the off-diagonal blocks of $\mathcal{A}_{\text{U}}$ implies that under this configuration, the contributions to the connection from the coupling between the ground and excited states vanish. To evaluate the Uhlmann phase, we also notice that the initial density matrix is given by
$\rho(0)=\frac{1}{Z}\mathrm{e}^{-\beta H(\theta(0),\phi(0))}$
$=$ $\left(\begin{array}[]{cccc}\frac{1}{4}&0&-\frac{\tanh(\beta R)}{4\sqrt{2}}&-\frac{\tanh(\beta R)}{4\sqrt{2}}\\ 0&\frac{1}{4}&-\frac{\tanh(\beta R)}{4\sqrt{2}}&\frac{\tanh(\beta R)}{4\sqrt{2}}\\ -\frac{\tanh(\beta R)}{4\sqrt{2}}&-\frac{\tanh(\beta R)}{4\sqrt{2}}&\frac{1}{4}&0\\ -\frac{\tanh(\beta R)}{4\sqrt{2}}&\frac{\tanh(\beta R)}{4\sqrt{2}}&0&\frac{1}{4}\\ \end{array}\right).$ (29)
Introducing $\chi\equiv 1-\operatorname{sech}(R/T)$, the Uhlmann holonomy is evaluated as
$\mathcal{P}\mathrm{e}^{-\oint_{C}\mathcal{A}_{\text{U}}}$ $=\mathcal{P}\exp\left(-\mathrm{i}\oint_{C}\frac{\chi}{2}\mathrm{d}\phi\sigma_{3}\oplus(-\sigma_{1})\right)$
$=\exp(-\mathrm{i}\pi\chi\sigma_{3})\oplus\exp(\mathrm{i}\pi\chi\sigma_{1})$
$=\begin{pmatrix}\mathrm{e}^{-\mathrm{i}\pi\chi}&0&0&0\\ 0&\mathrm{e}^{\mathrm{i}\pi\chi}&0&0\\ 0&0&\cos(\pi\chi)&\mathrm{i}\sin(\pi\chi)\\ 0&0&\mathrm{i}\sin(\pi\chi)&\cos(\pi\chi)\end{pmatrix},$ (30)
Using Eq. (III.1), the corresponding Uhlmann phase is then given by
$\theta_{\text{U}}(C)=\arg\text{Tr}\left[\rho(0)\mathcal{P}\mathrm{e}^{-\oint_{C}\mathcal{A}_{\text{U}}}\right]=\arg\left(\cos(\pi\chi)\right).$ (31)
For the WZ connection in the degenerate ground-state subspace, it can be shown that $\mathcal{A}_{\pm}=\frac{1\pm\cos\theta}{2\mathrm{i}}\mathrm{d}\phi\sigma_{1}$ for this simple case. If the same path along the equator is followed with $\theta=\frac{\pi}{2}$, the WZ connection of the degenerate ground states reduces to
$\mathcal{A}_{-}=-\frac{\mathrm{i}}{2}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\mathrm{d}\phi.$ (32)
Again, the path-ordering can be omitted since the connection is proportional to a constant matrix. The corresponding holonomy of the degenerate ground states is given by
$U_{-}=\mathrm{e}^{\frac{\mathrm{i}}{2}\sigma_{1}{\oint}\mathrm{d}\phi}=-1_{2}.$ (33)
Using Eq. (21), we find that
$\theta_{\text{WZ}}=\arg\text{Tr}_{-}\left(-\frac{1}{2}P_{-}\right)=\arg(-1)=\pi.$ (34)
As a comparison, when $T\to 0$, we have $\beta\to\infty$ and $\chi\to 1$. Thus, Eq. (31) implies
$\lim_{T\to 0}\theta_{\text{U}}=\arg(\cos(\pi))=\pi=\theta_{\text{WZ}}.$ (35)
Therefore, this simple case shows a correspondence between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase.
Figure 1: Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ for the simple case. A $\pi$-jump occurs at $T\approx 0.75R$.
Before proceeding further, we note that the behavior of the Uhlmann phase often signals finite-temperature topological phase transitions [31, 34, 35]. At infinitely high temperatures, the density matrix becomes proportional to the identity matrix, implying $\rho\propto 1_{4}$ for a four-level system. As a result, any evolution of $\rho(t)$ along a loop in the parameter space effectively collapses to a single point, thereby rendering the Uhlmann phase trivial [34]. Since $\lim_{T\to 0}\theta_{\text{U}}=\pi$, there must exist a critical temperature $T_{c}$ at which the Uhlmann phase undergoes a sudden jump, indicating a topological phase transition. For the simple case studied here, the critical temperature showing a jump of the Uhlmann phase emerges when $\cos(\pi\chi)=0$ or $\chi=1/2$ since $10<\chi<$ 1 at finite temperatures, which yields $T_{c}\approx 0.75R$. To visualize the transition, we plot $\theta_{\text{U}}$ as a function of $T$ on a semi-logarithmic scale in Figure 1. As shown in the plot, the Uhlmann phase indeed jumps by a quantized value of $\pi$. The change in the Uhlmann phase indicates that below $T_{c}$, the final purification $W(\tau)$ is antiparallel to the initial purification $W(0)$ while above $T_{c}$, $W(\tau)$ becomes parallel to $W(0)$.
III.2 4D tight-binding model
Next, we consider a more complex case of a four-band model that realizes a four-dimensional (4D) tight-binding Hamiltonian, originally introduced in Ref. [43]. The Hamiltonian in real space is a lattice model given by
$H=\sum_{n}\left[\sum_{j=1}^{4}\left(\psi^{\dagger}_{n}\frac{\Gamma^{0}-\mathrm{i}\Gamma^{j}}{2}\psi_{n+j}+\text{H.c.}\right)+m\psi^{\dagger}_{n}\Gamma^{0}\psi_{n}\right].$ (36)
Here $n$ and $j$ denote the lattice location and its four nearest-neighbors. The real-space Hamiltonian can be transformed into the following expression in momentum space
$H=\sum_{\mathbf{k}}\psi^{\dagger}_{\mathbf{k}}R_{a}\Gamma^{a}\psi_{\mathbf{k}}.$ (37)
This corresponds to the four-level model of Eq. (1) by identifying the components of the vector $\mathbf{R}$ as $R_{1}=\sin k_{x}$, $R_{2}=\sin k_{y}$, $R_{3}=\sin k_{z}$, $R_{4}=\sin k_{u}$, and $R_{5}=m+\cos k_{x}+\cos k_{y}+\cos k_{z}+\cos k_{u}$. The parameter space is the first Brillouin zone with $k_{x}$, $k_{y}$, $k_{z}$, and $k_{u}$ spanning the interval $[-\pi,\pi]$, which forms a four-dimensional torus $T^{4}$.
For the 4D tight-binding model, the full expression of the Uhlmann connection is provided in Appendix A. We consider a special loop in $T^{4}$ defined by $C(t):=(k_{x}(t),k_{y}(t),k_{z}(t),k_{u}(t))=(k_{x}(t),0,0,0)$ and evolve the system accordingly. In this case, the Uhlmann connection reduces to
$\mathcal{A}_{\text{U}}=$ $-\frac{1-\operatorname{sech}(\beta R)}{2R^{2}}\begin{pmatrix}0&0&0&-1\\ 0&0&-1&0\\ 0&1&0&0\\ 1&0&0&0\end{pmatrix}$
×$\times\left[(m+3)\cos k_{x}+1\right]\mathrm{d}k_{x},$ (38)
To evaluate the Uhlmann phase, we define
$I(m,T):={\int}_{0}^{2\pi}\frac{\operatorname{sech}(\beta R)-1}{2R^{2}}\left[(m+3)\cos k_{x}+1\right]\mathrm{d}k_{x},$ (39)
where $R=\sqrt{(m+3)^{2}+2(m+3)\cos k_{x}+1}$. The Uhlmann holonomy in this case is given by
$\mathcal{P}\mathrm{e}^{-\oint_{C}\mathcal{A}_{\text{U}}}=\left(\begin{array}[]{cccc}\cos(I)&0&0&\sin(I)\\ 0&\cos(I)&\sin(I)&0\\ 0&-\sin(I)&\cos(I)&0\\ -\sin(I)&0&0&\cos(I)\\ \end{array}\right).$ (44)
Suppose the evolution starts at $k_{x}=0$, then the initial density matrix is
$\rho(0)=\frac{1}{4}(1_{4}-\tanh(\beta R)\Pi).$ (45)
Here $\Pi=\operatorname{diag}(1,1,-1,-1)$. With the setup, we can numerically evaluate the Uhlmann phase given by
$\theta_{\text{U}}=\arg\operatorname{Tr}[\rho(0)\mathrm{e}^{-\oint_{C}\mathcal{A}_{\text{U}}}]=\arg(\cos(I(m,T))).$ (46)
Figure 2: (Top panel) Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for $m=-3$. A $\pi$-jump occurs at $T\approx 0.75R_{0}$, signaling a topological phase transition. (Bottom panel) Uhlmann-phase diagram of the 4D tight-binding model. The numerical values of $T_{c}$ are marked by the blue circles. The background and dome-shape regions correspond to the trivial ($\theta_{\text{U}}=0$) and topological ($\theta_{\text{U}}=\pi$) phases, respectively. The scalar WZ phase at $T=0$, however, vanishes after the system traverses the path $C$ regardless of $m$. Here $R_{0}=R(m=-3)$.
In the top panel of Fig. 2, we plot the Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for $m=-3$. A clear $\pi$-jump is observed at the critical temperature $T_{c}=0.75R_{0}$, where $R_{0}=R(m=-3)$. In the bottom panel, we present the Uhlmann-phase diagram of the 4D tight-binding model on the $m\text{-}T$ plane. A finite Uhlmann phase appears exclusively in the range $-4<m<-2$ at $T=0$ and forms a dome-shape region at finite temperatures. The rest of the diagram is occupied by the topologically trivial region with zero Uhlmann phase. By fitting the boundary of the dome-shaped topological region with a symmetric power-law envelope function, we obtain an approximate relation between $T_{c}$ and $m$: $T_{c}/R_{0}\approx 0.75\left[1-(m+3)^{2}\right]^{0.45}$.
For the calculation of the scalar WZ phase of the ground states of the 4D tight-binding model, we use Eqs. (II.3) and (19) to find
$\mathcal{A}_{-}=0$ (47)
in this case. Consequently, the scalar WZ phase vanishes after the loop $C$ is traversed in the parameter space, regardless of the value of $m$. This explicit example demonstrates that the $T\rightarrow 0$ limit of $\theta_{\text{U}}$ does not always reduce to $\theta_{\text{WZ}}$, showcasing a fundamental difference from the previous case. As explained below Eq. (22), the scalar WZ phase is a valid geometric phase of the degenerate ground states, not given simply by the $T\rightarrow 0$ limit of the Uhlmann phase.
IV Correspondence between the Uhlmann and scalar WZ phases
The two examples sharply contrast different outcomes regarding the relationships between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase. We now turn to a more general analysis of this issue from an analytic perspective. Our objectives are twofold: to determine whether a simple relationship exists between the Uhlmann connection in the zero-temperature limit and the WZ connection and to identify the conditions under which the Uhlmann phase in the zero-temperature limit may reduce to the scalar WZ phase. In the following, we consider generic systems with $N$-dimensional Hilbert space.
In a system where the $i$-th eigenstate $|\psi^{(i)}_{n}\rangle$ has degeneracy $D_{i}$, with $n=1,\dots,D_{i}$ and $\sum_{i}D_{i}=N$, the Uhlmann connection $\mathcal{A}_{\text{U}}$ in Eq. (20) can be more precisely formulated as:
$\mathcal{A}_{\text{U}}=$ $-\sum_{i\neq j}\sum_{n=1}^{D_{i}}\sum_{m=1}^{D_{j}}\frac{\left(\sqrt{\lambda_{i}}-\sqrt{\lambda_{j}}\right)^{2}}{\lambda_{i}+\lambda_{j}}$
×$\times|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|,$ (48)
where $|\mathrm{d}\psi^{(j)}_{m}\rangle\equiv\mathrm{d}|\psi^{(j)}_{m}\rangle$. Here $\lambda_{n}=\frac{e^{-\beta E_{n}}}{Z}$ with the energy levels ordered as $E_{0}<E_{1}<\dots$. Since $\lambda_{i}\neq\lambda_{j}$, we define $\lambda_{m}=\min\{\lambda_{i},\lambda_{j}\}$ and $\lambda_{M}=\max\{\lambda_{i},\lambda_{j}\}$. In the low-temperature limit, we find,
$\lim_{T\to 0}\frac{\left(\sqrt{\lambda_{i}}-\sqrt{\lambda_{j}}\right)^{2}}{\lambda_{i}+\lambda_{j}}=\lim_{T\to 0}\frac{\left(1-\sqrt{\frac{\lambda_{m}}{\lambda_{M}}}\right)^{2}}{1+\frac{\lambda_{m}}{\lambda_{M}}}=1.$ (49)
When $T\to 0$, the Uhlmann connection can be expressed as
$\mathcal{A}_{\text{U}}\to$ $-\sum_{i\neq j}\sum_{n=1}^{D_{i}}\sum_{m=1}^{D_{j}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|$
$=$ $-\sum_{i,j}\sum_{n=1}^{D_{i}}\sum_{m=1}^{D_{j}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|$
$+\sum_{i}\sum_{n,m=1}^{D_{i}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(i)}_{m}\rangle\langle\psi^{(i)}_{m}|$
$=$ $-\sum_{j,m}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|+\mathcal{A}_{\text{WZ}},$ (50)
where
$\mathcal{A}_{\text{WZ}}=$ $\sum_{i}\sum_{n,m=1}^{D_{i}}\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(i)}_{m}\rangle|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{m}|$
$=$ $\begin{pmatrix}\mathcal{A}^{(0)}_{\text{WZ}}&&\\ &\mathcal{A}^{(1)}_{\text{WZ}}&\\ &&\ddots\end{pmatrix}$ (51)
is the full WZ connection matrix. Here we have suppressed the explicit dependence on the parameter $\mathbf{R}$ for brevity. According to Eq. (IV), the Uhlmann holonomy along a closed loop $C$ in the parameter space is expressed as:
$\lim_{T\to 0}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}=\mathcal{P}\mathrm{e}^{-{\oint}_{C}\left(\mathcal{A}_{\text{WZ}}-\sum_{j,m}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|\right)}.$ (52)
In the first example above, when the system evolves along the equator, the eigenvalues ±$\pm R$ remain constant, and the associated evolution corresponds to a unitary transformation of the states involved. This can be written as $|\psi^{(j)}_{m}(\mathbf{R}(t))\rangle=\mathcal{D}(\mathbf{R}(t))|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. For notational simplicity, we denote $|\psi^{(j)}_{m}(0)\rangle=|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. In this scenario, the relationship between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ simplifies to
$\mathcal{A}_{\text{U}}\to$ $\mathcal{A}_{\text{WZ}}-\sum_{j,m}\mathrm{d}\mathcal{D}|\psi^{(j)}_{m}(0)\rangle\langle\psi^{(j)}_{m}(0)|\mathcal{D}^{\dagger}$
$=\mathcal{A}_{\text{WZ}}-\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}.$ (53)
Thus, we obtain a direct relation that explicit connects the Uhlmann connection and the Wilczek-Zee connection. Moreover, the relation (52) simplifies to
$\lim_{T\to 0}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}=\mathcal{P}\mathrm{e}^{-{\oint}_{C}\left(\mathcal{A}_{\text{WZ}}-\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}\right)}.$ (54)
Assuming the additional condition $[\mathcal{A}_{\text{WZ}},\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}]=0$ holds, the expression further reduces to
$\lim_{T\to 0}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}=\left(\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{WZ}}}\right)\left(\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}\right).$ (55)
Since $\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}$ is anti-Hermitian, the path-ordered exponentiation $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ forms an $N\times N$ unitary matrix. Given that this path-ordered exponentiation maps the closed loop $C\sim S^{1}$ to the group $U(N)$ of unitary matrices, the resulting $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ is an element of the fundamental group $\pi_{1}(U(N))=\mathbb{Z}$. Thus, we obtain [44]
$\mathcal{P}\exp\left({\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}\right)=\mathrm{e}^{2\pi\mathrm{i}\kappa}1_{N},$ (56)
where $1_{N}$ denotes the $N\times N$ identity matrix, and the winding number $\kappa$ associated with $\mathcal{D}$ along the loop $C$ is given by:
$\kappa=\frac{1}{2\pi\mathrm{i}}\oint_{C}\text{Tr}\left(\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}\right).$ (57)
If $C$ is a loop contractible to a point, then $\kappa=0$. In this case,
$\lim_{T\to 0}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{U}}}=\mathrm{e}^{2\pi\mathrm{i}k}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{\text{WZ}}}.$ (58)
Moreover, in the low temperature limit,
$\rho\to\frac{1}{D_{0}}\sum_{n}|\psi^{(0)}_{n}\rangle\langle\psi^{(0)}_{n}|=\begin{pmatrix}\frac{1}{D_{0}}P_{0}&&\\ &0&\\ &&\ddots,\end{pmatrix},$ (59)
where $P_{0}$ is the projector to the ground state with degeneracy $D_{0}$. The $T\rightarrow 0$ Uhlmann phase thus becomes
$\theta_{\text{U}}(C)$ $\to\arg\text{Tr}_{-}\left(\frac{1}{D_{0}}P_{0}\mathrm{e}^{2\pi\mathrm{i}\kappa}\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}^{(0)}_{\text{WZ}}}\right)$
$=\theta_{\text{WZ}}(C)+2\pi\kappa$
$=\theta_{\text{WZ}}(C)\mod 2\pi.$ (60)
The above argument establishes a conditional relation between the Uhlmann phase as $T\to 0$ and the scalar WZ phase when the system evolves unitarily under a transformation $\mathcal{D}$, provided the condition $[\mathcal{A}_{\text{WZ}},\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}]=0$ holds. However, this condition is somewhat too restricted. In fact, the first example we presented above (the simple case) does not satisfy this commutation relation, yet the correspondence between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase remains valid. This suggests that the actual requirement for the correspondence to hold may be more lenient, though a rigorous mathematical characterization still awaits future research. In contrast, the second example of the 4D tight-binding model shows non-unitary evolution since the eigenvalues vary as the system evolves in the parameter space. Therefore, the relationship between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase becomes considerably more subtle and dependent on specific features of the system. Our results thus illustrate the complex physics facing finite-temperature topological quantum systems.
V Implications
Degenerate systems have been of interest in holonomic quantum computation (HQC) [45, 46, 47], where the phase factors are engineered to achieve quantum gate operations. Some exemplary systems considered in this direction include Majorana fermions and atoms interacting with laser light. The WZ connection contributes to the non-Abelian holonomy, which is frequently used in various schemes of HQC. The introduction of the scalar WZ phase will help further quantify the holonomy induced by the WZ connection int the degenerate ground-state subspace. Moreover, quantum geometries of degenerate pure states have been characterized [48, 49, 50], which may inspire future research to bridge local geometry and global topology of degenerate quantum systems.
Since a typical Uhlmann process is incompatible with the Hamiltonian dynamics [51], measuring the Uhlmann phase in natural systems typically encounter the uncontrollable problem of the environment to satisfy the Uhlmann parallel-transport condition. Nevertheless, the Uhlmann phase of a two-level system has been simulated on the IBM quantum platform [52] because the ancilla in the simulation modeling the environment can be fully controlled to respect the Uhlmann parallel-transport condition. Further, there is a proposal for generalizing the scheme to spin-$j$ systems [34]. The four-level model analyzed in this work may be encoded by two qubits, which will require another two qubits to serve as the ancilla to enact the system-environment coupling for generating the Uhlmann phase. Although simulating the Uhlmann phase of degenerate systems on quantum computers can be a challenging task, its success will allow us to explore the rich physics behind finite-temperature topological systems.
The comparison between the $T\rightarrow 0$ results from the Uhlmann connection and those from the Berry or WZ connection of the corresponding pure states helps address the fundamental issue that although the density matrix smoothly transit from the $T\rightarrow 0$ mixed state to the $T=0$ pure state, the Uhlmann fiber bundle is for full-rank density matrices but the pure-state bundles only concerns the ground-state subspace. Therefore, any correspondence between the $T\rightarrow 0$ results and the pure-state results needs to be analyzed on a case-by-case basis [41]. The conditional correspondence between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase derived here may serve as a starting point for developing a broader classification of the correspondence between the two phases in degenerate quantum systems.
We mention that alternative approaches to mixed-state geometric phases have been proposed, including the interferometric geometric phase (IGP) [53, 54, 55, 56, 57, 58, 59] inspired by optical interferometry. The IGP provides an experimentally accessible phase observable without requiring cyclic evolution. Moreover, it can be shown that the IGP is equivalent to a weighted sum of the Berry phases over the eigenstates of the system for cyclic processes [53, 60]. The experimental realizations of the IGP in platforms such as NMR and polarized neutron systems [61, 62, 63, 64] underscore its practical value. There have been some attempts to generalize the IGP to degenerate systems [65]. However, the underlying parallel-transport condition of the IGP still awaits a concise fiber-bundle foundation to characterize its fully geometric meaning [66].
VI conclusion
We have investigated the mixed-state Uhlmann phase of an exemplary degenerate system and searched for possible relationships with the WZ holonomy in the degenerate ground-state subspace through the introduction of the scalar WZ phase. The four-level model featuring two doubly degenerate subspaces allows for explicit forms of the Uhlmann and WZ connections to evaluate their holonomy and phases. Through explicit examples, we show that the Uhlmann phase in the zero-temperature limit may or may not approach the scalar WZ phase of the degenerate ground states. Moreover, our theoretical analysis gives some conditions under which the agreement between the zero-temperature limit of the Uhlmann phase and the scalar WZ phase is guaranteed.
The rich phenomena behind the mixed-state Uhlmann phase and pure-state scalar WZ phases of systems with degeneracy highlight challenges facing geometric and topological properties of degenerate quantum systems at and beyond zero temperature. Our examples and analyses make some progress towards understanding the deeper connection between mixed-state and pure-state topology in the presence of degeneracy and offer inspiration for investigating temperature effects in topological systems with potential applications in condensed matter physics and quantum information science.
Acknowledgements.
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). C.C.C. was supported by the NSF (No. PHY-2310656) and DOE (No. DE-SC0025809).
Appendix A More details of the Uhlmann connection
Using the diagonal form $\sqrt{\rho}=\sum_{i}\sqrt{\lambda_{i}}|i\rangle\langle i|$, we obtain
$[\sqrt{\rho},\mathrm{d}\sqrt{\rho}]=$ $\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}\left(|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|-|j\rangle(\mathrm{d}\langle j|)|i\rangle\langle i|\right)$
$+$ $\sum_{i}\lambda_{i}\left(|i\rangle\mathrm{d}\langle i|-\mathrm{d}|i\rangle\langle i|\right).$ (61)
By exchanging indices $i\leftrightarrow j$ in the second term of the first line on the right-hand side and using $(\mathrm{d}\langle i|)|j\rangle=-\langle i|\mathrm{d}|j\rangle$, the first line becomes
$2\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|.$ (62)
Following a similar derivation, the second line becomes
$\sum_{i,j}\lambda_{i}\left(|i\rangle(\mathrm{d}\langle i|)|j\rangle\langle j|-|j\rangle\langle j|\mathrm{d}|i\rangle\langle i|\right)$
$=$ $-\sum_{i,j}(\lambda_{i}+\lambda_{j})|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|.$ (63)
Combining the above results, we obtain
$[\sqrt{\rho},\mathrm{d}\sqrt{\rho}]=-\sum_{i,j}\left(\sqrt{\lambda_{i}}-\sqrt{\lambda_{j}}\right)^{2}|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|.$ (64)
Substituting this into Eq. (9), we obtain an alternative expression for the Uhlmann connection:
$\mathcal{A}_{\text{U}}=$ $-\sum_{i\neq j}\frac{\left(\sqrt{\lambda_{i}}-\sqrt{\lambda_{j}}\right)^{2}}{\lambda_{i}+\lambda_{j}}|i\rangle\langle i|\mathrm{d}|j\rangle\langle j|$
$=$ $-\sum_{i\neq j}\left(1-\frac{2\sqrt{\lambda_{i}\lambda_{j}}}{\lambda_{i}+\lambda_{j}}\right)|i\rangle\langle i|\mathrm{d}\left(|j\rangle\langle j|\right),$ (65)
where the second line comes from the fact that $\langle i|j\rangle=0$ if $\lambda_{i}\neq\lambda_{j}$ in thermal equilibrium states. For the four-level model with two doubly degenerate spaces considered in the main text, $\lambda_{\pm}=\frac{\mathrm{e}^{\mp\beta R}}{4\cosh(\beta R)}$, satisfying $\lambda_{+}+\lambda_{-}=\frac{1}{2}$ and $\sqrt{\lambda_{+}\lambda_{-}}=\frac{1}{4\cosh(\beta R)}$. Substituting the above expressions into Eq. (A) and plugging in $|i\rangle,|j\rangle=\{|\psi_{a}\rangle,|\psi_{b}\rangle,|\psi_{c}\rangle,|\psi_{d}\rangle\}$, we get
$\mathcal{A}_{\text{U}}=-\left[1-\text{sech}\left(\frac{R}{T}\right)\right]\left(P_{+}\mathrm{d}P_{-}+P_{-}\mathrm{d}P_{+}\right),$ (66)
which then leads to Eq. (II.2).
For the 4D tight-binding model, the Uhlmann connection according to Eq. (II.2) is give by
$\mathcal{A}_{\text{U}}=-\frac{1-\operatorname{sech}(\beta R)}{2R^{2}}M,$ (67)
where the elements of the matrix $M$ are:
$\begin{array}[]{rcl}M_{11}&=&\mathrm{i}\left[\sin k_{y}\cos k_{x}\mathrm{d}k_{x}-\sin k_{x}\cos k_{y}\mathrm{d}k_{y}+\sin k_{u}\cos k_{z}\mathrm{d}k_{z}-\sin k_{z}\cos k_{u}\mathrm{d}k_{u}\right],\\ M_{12}&=&(\sin k_{x}-\mathrm{i}\sin k_{y})\left[\cos k_{z}\mathrm{d}k_{z}-\mathrm{i}\cos k_{u}\mathrm{d}k_{u}\right]-(\sin k_{z}-\mathrm{i}\sin k_{u})\left[\cos k_{x}\mathrm{d}k_{x}-\mathrm{i}\cos k_{y}\mathrm{d}k_{y}\right],\\ M_{13}&=&-\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{z}\mathrm{d}k_{z}+\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{u}\mathrm{d}k_{u}\\ &&\quad+(\sin k_{z}-\mathrm{i}\sin k_{u})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{14}&=&-\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{x}\mathrm{d}k_{x}+\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{y}\mathrm{d}k_{y}\\ &&\quad+(\sin k_{x}-\mathrm{i}\sin k_{y})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{21}&=&(\sin k_{z}+\mathrm{i}\sin k_{u})\left[\cos k_{x}\mathrm{d}k_{x}+\mathrm{i}\cos k_{y}\mathrm{d}k_{y}\right]-(\sin k_{x}+\mathrm{i}\sin k_{y})\left[\cos k_{z}\mathrm{d}k_{z}+\mathrm{i}\cos k_{u}\mathrm{d}k_{u}\right],\\ M_{22}&=&\mathrm{i}\left[-\sin k_{y}\cos k_{x}\mathrm{d}k_{x}+\sin k_{x}\cos k_{y}\mathrm{d}k_{y}-\sin k_{u}\cos k_{z}\mathrm{d}k_{z}+\sin k_{z}\cos k_{u}\mathrm{d}k_{u}\right],\\ M_{23}&=&(\sin k_{x}+\mathrm{i}\sin k_{y})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right]\\ &&\quad-\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\left[\cos k_{x}\mathrm{d}k_{x}+\mathrm{i}\cos k_{y}\mathrm{d}k_{y}\right],\\ M_{24}&=&\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{z}\mathrm{d}k_{z}+\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{u}\mathrm{d}k_{u}\\ &&\quad-(\sin k_{z}+\mathrm{i}\sin k_{u})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{31}&=&M_{24},\\ M_{32}&=&\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{x}\mathrm{d}k_{x}-\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{y}\mathrm{d}k_{y}\\ &&\quad-(\sin k_{x}-\mathrm{i}\sin k_{y})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{33}&=&\mathrm{i}\left[\sin k_{y}\cos k_{x}\mathrm{d}k_{x}-\sin k_{x}\cos k_{y}\mathrm{d}k_{y}-\sin k_{u}\cos k_{z}\mathrm{d}k_{z}+\sin k_{z}\cos k_{u}\mathrm{d}k_{u}\right],\\ M_{34}&=&(\sin k_{x}-\mathrm{i}\sin k_{y})\left[\cos k_{z}\mathrm{d}k_{z}+\mathrm{i}\cos k_{u}\mathrm{d}k_{u}\right]-(\sin k_{z}+\mathrm{i}\sin k_{u})\left[\cos k_{x}\mathrm{d}k_{x}-\mathrm{i}\cos k_{y}\mathrm{d}k_{y}\right],\\ M_{41}&=&\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{x}\mathrm{d}k_{x}+\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{y}\mathrm{d}k_{y}\\ &&\quad-(\sin k_{x}+\mathrm{i}\sin k_{y})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{42}&=&-\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{z}\mathrm{d}k_{z}+\mathrm{i}\left(m+\sum_{\alpha}\cos k_{\alpha}\right)\cos k_{u}\mathrm{d}k_{u}\\ &&\quad+(\sin k_{z}-\mathrm{i}\sin k_{u})\left[-\sin k_{x}\mathrm{d}k_{x}-\sin k_{y}\mathrm{d}k_{y}-\sin k_{z}\mathrm{d}k_{z}-\sin k_{u}\mathrm{d}k_{u}\right],\\ M_{43}&=&(\sin k_{z}-\mathrm{i}\sin k_{u})\left[\cos k_{x}\mathrm{d}k_{x}+\mathrm{i}\cos k_{y}\mathrm{d}k_{y}\right]-(\sin k_{x}+\mathrm{i}\sin k_{y})\left[\cos k_{z}\mathrm{d}k_{z}-\mathrm{i}\cos k_{u}\mathrm{d}k_{u}\right],\\ M_{44}&=&\mathrm{i}\left[-\sin k_{y}\cos k_{x}\mathrm{d}k_{x}+\sin k_{x}\cos k_{y}\mathrm{d}k_{y}+\sin k_{u}\cos k_{z}\mathrm{d}k_{z}-\sin k_{z}\cos k_{u}\mathrm{d}k_{u}\right].\end{array}$ (91)